Without actually performing the long division, state whether the following rational numbers will have a terminating decimal expansion or a non-terminating repeating decimal expansion:
23/(2^3 5^2 ) 

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

23/ (2^3 5^2)= 23/(8×25)= 23/200
23/ (2^3 5^2) = 0.115

Thanks